Perry's Pizzeria issued 20-year annual coupon bonds one year ago at a coupon rate of 5.1 percent. If the YTM on these bonds is 8.99 percent, what is the current bond price?
Answer #1

Information provided:

Face value= future value= $1,000

Time= 20 years - 1 year= 19 years

Coupon rate= 5.1%

Coupon payment= 0.051*1,000= $51

Yield to maturity= 8.99%

The current price of the bond is calculated by computing the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

PMT= 51

I/Y= 8.99

N= 19

Press the CPT key and PV to compute the present value.

The value obtained is 651.60.   

Therefore, the current price of the bond is $651.60.

In case of any query, kindly comment on the solution.